OpenStudy (anonymous):

5^-2 find an equivalent expression. write your answer as a fraction

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

take it to denominator

OpenStudy (mayankdevnani):

ALWAYS REMEMBER :- \[\large \bf x^{-1}=\frac{1}{x^1}\]

OpenStudy (gorv):

powwer will become positive there

OpenStudy (mayankdevnani):

\[\large \bf x^{-2}=\frac{1}{x^2}\] and so on

OpenStudy (anonymous):

so 1/25?

OpenStudy (anonymous):

or 25^2

OpenStudy (mayankdevnani):

correct :) its 1/25

OpenStudy (mayankdevnani):

because :- \[\large \bf 5^{-2}=\frac{1}{5^2}=\frac{1}{25}=\color{red}{Answer}\]

OpenStudy (mayankdevnani):

any questions ?? @mcnabkt

OpenStudy (anonymous):

thanks :) and no i understand it now

OpenStudy (anonymous):

i might later on though

OpenStudy (mayankdevnani):

good :)